# Riddle Me This! Stealthy Membership Inference for Retrieval-Augmented Generation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06copyright: nonecopyright: nonefootnotetext: Equal Contributionfootnotetext: Correspondence to anaseh@cs.umass.edu
@@ 1. Introduction
Membership inference attacks (MIAs) in machine learning attempt to discern if a given record was part of a given model’s training data. MIAs thus have great utility for privacy auditing, copyright violations (Maini et al., 2024), and test-set contamination (Oren et al., 2023). While MIAs generally relate to the information contained in the model’s parameters (with the model having seen some data during training), inferring the presence of particular documents in a RAG’s data-store is different as the knowledge is not directly contained in model parameters. Inferring the membership of such documents can directly damage one of the primary benefits of RAG for privacy. With private knowledge being non-parametric, existing membership-inference attacks that rely on the memorization of training data may not work.
@@
Towards this, we systematically evaluate existing MIAs (Liu et al., 2024b; Anderson et al., 2024; Li et al., 2024b) across various detection mechanisms and show that prior attacks completely break down against these detection strategies (Section 4). We then introduce Interrogation Attack (IA), a MIA which is:
@@
Inspired by the doc2query task in Information Retrieval (IR) literature (Nogueira et al., 2019; Gospodinov et al., 2023), we employ established few-shot prompting (Dai et al., 2023) techniques to guide a LLM in creating queries that are both topically aligned with and uniquely answerable by the target document. These queries capture fine-grained and nuanced information specific to the target document, enabling us to subtly exploit the behavior of the RAG system in an undetectable manner. We then issue these queries to the RAG system. Since they are highly relevant to the target document ($d^{*}$), a well-performing RAG will retrieve and incorporate $d^{*}$ (if available) to generate accurate answers. We can thus verify the correctness of these answers to probe membership. Aggregating signals from multiple queries enables strong membership inference. Crucially, each query remains benign, avoiding direct requests for verbatim content or displaying suspicious “jailbreaking” patterns, ensuring the attack remains undetectable by any detection systems.
We conduct extensive experiments across multiple datasets and RAG configurations by varying retrieval and generation models (Section 6). While existing attacks are either detected easily or lack potency, we achieve successful inference while remaining virtually indistinguishable from natural queries, with detection rates as low as 5%, compared to upwards of 90% for most inference attacks against RAG. Finally, we analyze our attack’s failure cases (Section 7) and find that RAG may often be unnecessary: in many instances, the underlying LLM can answer questions about a given document without direct access to it, thereby questioning the necessity of a RAG-based system for such scenarios.
@@ 2.1. Retrieval Augmented Generation (RAG)
where $\text{ins}(q,D_{k})$ represents the query and context wrapped in a system instruction for the generative model An end user only gets to submit query $q$ to the RAG system $\mathcal{S}$ and observe the response $y$ directly in the form of generated text.
@@ 2.3. Privacy Attacks in RAGs
Recent research has explored various inference attacks against RAG systems. Anderson et al. (2024) developed techniques across different access levels, including a gray-box method using a meta-classifier on model logits and a black-box approach directly querying model membership. Li et al. (2024b) a similarly straightforward approach, where the target document is broken into two parts, with the idea that presence of the target document in the context would lead the LLM into completing the given query (one half of the document). However, authors for both these works find that simple modifications to the system instruction can reduce attack performance significantly to near-random.
@@
Several recent works have also proposed context leakage and integrity attacks, where the adversary has the capability of inject malicious documents into RAG knowledge database (Chaudhari et al., 2024; Jiang et al., 2024) or can poison the RAG system direcly (Peng et al., 2024). This threat model is different than ours as we do not assume any RAG poisoning or knowledge base contamination for our MIA.
@@ 3. Threat Model
$\mathcal{A}(d^{*})=\begin{cases}1,&\text{if }d^{*}\in\mathcal{D}\\ 0,&\text{if }d^{*}\notin\mathcal{D}\end{cases}$
@@
  • PII-Containing Documents: Documents with sensitive details, such as addresses or health records, present a high risk. Inferring their presence could result in severe privacy violations and potential regulatory breaches.
  • Factual Knowledge Sources: Documents containing specialized knowledge, such as internal manuals, proprietary research, or compliance guidelines, are often harder to target due to overlapping information across multiple documents. However, a successful inference in such cases could compromise intellectual property or reveal sensitive strategic information.
@@ 4. Limitations of Existing Inference Attacks on RAG Systems
Table 1. Evaluating the effectiveness of malicious-input detection against prompt-based methods proposed in the literature for privacy leakage (via membership inference or dataset extraction) for RAG-based systems. For prompts that require the query, we compute scores based on aggregate statistics over 500 samples from various datasets. Both few-shot GPT-4 and Lakera can easily detect attempts to infer retrieved documents. Our attack achieves near-zero detection rate, unlike prior attacks that are almost always detected.
@@
To evaluate the ability of current attacks to bypass detection methods, we adopt two different approaches. First we utilize LakeraGuard, a commercial off-the-shelf guard model designed for detecting prompt-injection and jailbreak attempts (Li et al., 2024a), to evaluate queries from different attacks. While this tool can detect queries from some existing attacks, it tends to fall short in identifying queries from attacks whose prompts appear more natural. These tools are designed to detect a wide range of prompt injection queries, so it is unsurprising that they may not perform perfectly in specialized settings like context probing attacks. To develop a more tailored detection tool, we leverage the capabilities of GPT-4o as a classifier with few-shot prompting to classify input queries as either ”natural” or ”context probing.”Instruction template for this classification task is presented in the Appendix (Figure 11) GPT-4o has recently shown great performance in prompt injection detection, further supporting its use for this task (Li et al., 2024a). Both approaches have shown good performance in prompt injection detection (Liu et al., 2024a).
@@
These results highlight the necessity for attackers to craft stealthy queries that avoid explicit instructions aimed at recovering documents from the model’s context. While adversarially crafted texts designed to bypass detection are feasible, an ideal attack strategy would involve generating clean-text queries that are immune to such defensive countermeasures. Thus for an inference attack to be successful in the context of a practical RAG system, it must bypass any query-filtering systems that can detect obvious inference attempts.
@@ 5.3. Membership Inference
(5) $S=\frac{1}{n}\sum_{i=1}^{n}\big{(}\mathbb{I}[r_{i}=g_{i}]-\lambda\mathbb{I}[r_{i}=\text{UNK}]\big{)},$
where $\mathbb{I}[\cdot]$ is the indicator function that evaluates to $1$ if the equality condition holds and 00 otherwise, and $\lambda$ is a hyper-parameter that penalizes the inability to answer a question. A higher score $S$ indicates that the RAG system consistently retrieves correct information, suggesting that $d^{*}$ is included in the knowledge base.
@@ 6.1. Evaluation Setup
Generator and Retriever. We utilize two retrievers in our evaluations: GTE (Li et al., 2023) and BGE (Li et al., 2023). For generators, we evaluate four different models: Llama 3.1 Instruct-8B (Dubey et al., 2024), Command-R-7Bhttps://huggingface.co/CohereForAI/c4ai-command-r7b-12-2024, Microsoft Phi-4 (Abdin et al., 2024), and Gemma-2-2B (Team et al., 2024).
@@
Query Generation Setting. For IA, we employ few-shot prompting with GPT-4o to generate 30 queries based on the target document. We also use GPT-4o to generate a short description of the target document, summarizing its main idea and keywords. For details of different prompting strategies and the corresponding prompts for each stage, see Appendix B and Appendix E.
RAG Setting. As described in Section 2.1, we evaluate our attack in a more realistic setting compared to previous works, where the RAG system employs query-rewriting on the user’s query. We implement query-rewriting using a simple query-paraphrasing prompt via GPT-4o. We set $k=3$ for retrieval and investigate the impact of this hyperparameter across all attacks in Section 6.3.2. These retrieved documents are then provided as context to the generator via a system prompt. Details on both the query-paraphrasing and system prompts are presented in Appendix E. To demonstrate the impact of query-rewriting on inference, we also evaluate attacks in a vanilla RAG setup where query-rewriting is disabled (Appendix D).
Figure 4. ROC curves for Command-R (7B) as generator, GTE as retriever, across various datasets. Our attack (IA) achieves near-perfect inference across multiple datasets. ROC curves for other RAG configurations, can be found in Appendix G.
Baselines. We compare our attack with three prior black-box MIAs against RAG systems: RAG-MIA (Anderson et al., 2024), $S^{2}$ (Li et al., 2024b), and MBA (Liu et al., 2024b). RAG-MIA takes a simpler approach by directly probing the RAG system to ask if the target document appears in the context. $S^{2}$ uses the first half of the target document as a query and calculates the semantic similarity score (i.e., BLEU) between the RAG system’s responses and the original document as the membership score. They hypothesize that if a document is present in the database, the RAG system’s responses will exhibit high semantic similarity to the original content. MBA uses a proxy-LM to selectively mask words in the target document, followed by querying the RAG to predict those masked words. The number of successfully predicted masked words is used as the membership score. In our experiments, we use Qwen-2.5-1.5B (Yang et al., 2024) as the proxy LM.
@@ 6.3. Ablation Study
Here we evaluate the impact of varying the number of queries made by our attack to the RAG (Section 6.3.1), as well as the impact of the number of documents retrieved as context by the RAG system (Section 6.3.2).
@@ 7.1. Assumptions on RAG Documents
Generic Documents. In addition to length, the targeted documents must be sufficiently informative. The attack may perform poorly on highly generic documents, as they do not contain enough specific details to craft unique and distinguishable questions. However, it is worth noting that in such cases, the utility of RAG might be limited, as generic documents provide less value for retrieval-based systems and the RAG owner might benefit in efficiency from discarding such documents from their datastore.
@@ 7.2. Analyzing Failure Cases
False Negatives. The fact that we observe high retrieval recall for our attack rules out the possibility of the target document being absent from the context provided to the RAG generator. The RAG’s inability to answer the question properly can thus have two potential reasons. On rare occasions, GPT-4o fails to paraphrase the user’s query accurately (see Appendix F for an example), which reflects a shortcoming in the RAG system—not being able to paraphrase a normal, benign query. For other cases, the RAG generator may struggle to answer the question even when the appropriate document is present in the provided context. Similarly, this can be attributed to the RAG’s generator lacking capabilities—especially given the fact that the question, by design, can be answered by GPT-4o-mini under the presence of the target document.
@@
False Positives. The RAG answering our queries correctly implies that the target document (corresponding to the query) is not required specifically as context to respond correctly. This can happen if similar documents with the necessary context are fetched by the retriever, or if the generator already possess sufficient knowledge to answer the question without relying on any context (see Appendix F for examples). To better understand this failure case, we compute the similarity between a non-member document $d$ and the document actually retrieved as context for a query corresponding to that non-member $d$, across multiple non-member documents and their corresponding queries generated for our attack. In Figure 8, we look at $n$-gram overlap and cosine similarity between retriever embeddings, and visualize them with respect to MIA scores for our attack. We observe that above some certain meaningful threshold ($0.2$ for 4-gram overlap, $0.9$ for embedding cosine similarity), there is a positive correlation between how ”similar” the non-member documents are to documents already present in the RAG, and the MIA Score (and by extension, questions answered correctly by the RAG). In summary, the failure cases are primarily due to limitations of the RAG system itself, such as occasional paraphrasing failures and the generator’s inability to answer questions effectively, rather than drawbacks of our attack.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3744840isbn: 979-8-4007-1525-9/2025/10ccs: Security and privacyccs: Computing methodologies Machine learningcopyright: nonefootnotetext: Equal Contributionfootnotetext: Correspondence to anaseh@cs.umass.edu
@@ 1. Introduction
Membership inference attacks (MIAs) in machine learning attempt to discern if a given record was part of a given model’s training data. MIAs thus have great utility for privacy auditing, copyright violations (Maini et al., 2024), and test-set contamination (Oren et al., 2023). While MIAs generally relate to the information contained in the model’s parameters (with the model having seen some data during training), inferring the presence of particular documents in a RAG’s data-store is different as the knowledge is not directly contained in model parameters. RAG systems introduce unique risks: even without exposing full content, simply confirming that a document is indexed can compromise privacy or reveal sensitive internal context. For example, documents containing PII might disclose that a user interacted with a system, while the presence of internal guidelines or strategy papers can hint at organizational priorities. These inferences carry implications for privacy, IP exposure, and regulatory compliance. Since RAG systems ingest data at deployment, membership inference can serve as a valuable tool for auditing what content—licensed, protected, or otherwise—has been incorporated.
@@
Towards this, we systematically evaluate existing MIAs (Liu et al., 2024b; Anderson et al., 2024; Li et al., 2024c) across various detection mechanisms and show that prior attacks completely break down against these detection strategies (Section 4). We then introduce Interrogation Attack (IA), a MIA which is:
@@
Inspired by the doc2query task in Information Retrieval (IR) literature (Nogueira et al., 2019; Gospodinov et al., 2023), we employ established few-shot prompting (Dai et al., 2023) techniques to guide an LLM in creating queries that are both topically aligned with and uniquely answerable by the target document. These queries capture fine-grained and nuanced information specific to the target document, enabling us to subtly exploit the behavior of the RAG system in an undetectable manner. We then issue these queries to the RAG system, with document $d^{*}$ being the target for membership inference. Since these queries are highly relevant to the target document, a well-performing RAG will retrieve and incorporate $d^{*}$ (if available) to generate accurate answers. We can thus verify the correctness of these answers to probe membership. Aggregating signals from multiple queries enables strong membership inference. Crucially, each query remains benign, avoiding direct requests for verbatim content or displaying suspicious “jailbreaking” patterns, ensuring the attack remains undetectable by any detection systems.
We conduct extensive experiments across multiple datasets and RAG configurations by varying retrieval and generation models (Section 6). While existing attacks are either detected easily or lack potency, we achieve successful inference while remaining virtually indistinguishable from natural queries, with detection rates as low as 5%, compared to upwards of 90% for most inference attacks against RAG. Finally, we analyze our attack’s failure cases (Section 7) and find that RAG may often be unnecessary: in many instances, the underlying LLM can answer questions about a given document without direct access to it, thereby questioning the necessity of a RAG-based system for such scenarios. Code for our experiments is available at https://github.com/ali7naseh/RAG_MIA.
@@ 2.1. Retrieval Augmented Generation (RAG)
where $\text{ins}(q,D_{k})$ represents the query and context wrapped in a system instruction for the generative model. An end user only gets to submit query $q$ to the RAG system $\mathcal{S}$ and observe the response $y$ directly in the form of generated text.
@@ 2.3. Privacy Attacks in RAGs
Recent research has explored various inference attacks against RAG systems. Anderson et al. (2024) developed techniques across different access levels, including a gray-box method using a meta-classifier on model logits and a black-box approach directly querying model membership. Li et al. (2024c) propose a similarly straightforward approach, where the target document is broken into two parts, with the idea that presence of the target document in the context would lead the LLM into completing the given query (one half of the document). However, authors for both these works find that simple modifications to the system instruction can reduce attack performance significantly to near-random.
@@
Several recent works have also proposed context leakage and integrity attacks, where the adversary has the capability of injecting malicious documents into RAG knowledge database (Chaudhari et al., 2024; Jiang et al., 2024) or can poison the RAG system direcly (Peng et al., 2024). This threat model is different than ours as we do not assume any RAG poisoning or knowledge base contamination for our MIA.
@@ 3. Threat Model
$\mathcal{A}(d^{*},\mathcal{S})=\begin{cases}1,&\text{if }d^{*}\in\mathcal{D}\\ 0,&\text{if }d^{*}\notin\mathcal{D}\end{cases}$
@@
  • PII-Containing Documents: Documents that contain personally identifiable information—such as internal user records, support tickets, financial transactions, or health-related forms—may not need to be leaked in full for privacy to be compromised. The mere confirmation that a particular document is part of the retrieval corpus can reveal that an individual engaged with a system, received a specific service, or appears in a sensitive internal context. Such inferences may already constitute privacy violations under data protection regulations like GDPR, particularly when tied to specific individuals.
  • Factual Knowledge Sources: Internal documentation such as policy guidelines, compliance manuals, proprietary research summaries, or strategic planning documents often contain overlapping factual content. While these documents may be more difficult to target directly, a successful membership inference can still reveal valuable information. For example, confirming that a particular regulatory checklist or internal strategy document is part of the RAG knowledge base may expose ongoing initiatives, compliance focus areas, or future product directions—information that can be strategically sensitive even in the absence of direct content leakage.
MIAs are therefore relevant not only from a privacy or IP exposure standpoint, but also for auditing and compliance purposes. Because RAG systems ingest documents at deployment time, the ability to determine document membership enables verification of whether sensitive, protected, or unlicensed content was indexed. This can assist with GDPR audits or copyright enforcement. At the same time, the same capability can be exploited as a preliminary step in more sophisticated data extraction attacks (Carlini et al., 2021).
@@ 4. Limitations of Existing Inference Attacks on RAG Systems
Table 1. Evaluating the effectiveness of malicious-input detection against prompt-based methods proposed in the literature for privacy leakage (via membership inference or dataset extraction) for RAG-based systems. Most attacks (except MBA) use the target sample directly in the attack query {sample}, as descbribed by the attack templates above. For prompts that require the query, we compute scores based on aggregate statistics over 500 samples from various datasets. Both few-shot GPT-4 and Lakera can easily detect attempts to infer retrieved documents. Our attack achieves near-zero detection rate, unlike prior attacks that are almost always detected.
@@
To evaluate the ability of current attacks to bypass detection methods, we adopt two different approaches. First we utilize LakeraGuard, a commercial off-the-shelf guard model designed for detecting prompt-injection and jailbreak attempts (Li et al., 2024b), to evaluate queries from different attacks. While this tool can detect queries from some existing attacks, it tends to fall short in identifying queries from attacks whose prompts appear more natural. These tools are designed to detect a wide range of prompt injection queries, so it is unsurprising that they may not perform perfectly in specialized settings like context probing attacks. To develop a more tailored detection tool, we leverage the capabilities of GPT-4o as a classifier with few-shot prompting to classify input queries as either ”natural” or ”context probing.”Instruction template for this classification task is presented in the Appendix (Figure 13) GPT-4o has recently shown great performance in prompt injection detection, further supporting its use for this task (Li et al., 2024b). Both approaches have shown good performance in prompt injection detection (Liu et al., 2024a).
@@
These results highlight the necessity for attackers to craft stealthy queries that avoid explicit instructions aimed at recovering documents from the model’s context. While adversarially crafted texts designed to bypass detection are feasible, an ideal attack strategy would involve generating clean-text queries that are immune to such defensive countermeasures. Making an attack stealthy requires it to be designed with stealth as a core objective. In Appendix E, we explore modifications to existing attacks to improve their stealth, but find that they remain largely detectable. Thus for an inference attack to be successful in the context of a practical RAG system, it must bypass any query-filtering systems that can detect obvious inference attempts.
@@ 5.3. Membership Inference
(5) $\frac{1}{n}\sum_{i=1}^{n}\big{(}\mathbb{I}[r_{i}=g_{i}]-\lambda\mathbb{I}[r_{i}=\text{UNK}]\big{)},$
where $\mathbb{I}[\cdot]$ is the indicator function that evaluates to $1$ if the equality condition holds and 00 otherwise, and $\lambda$ is a hyper-parameter that penalizes the inability to answer a question. A higher score indicates that the RAG system consistently retrieves correct information, suggesting that $d^{*}$ is included in the knowledge base.
@@ 6.1. Evaluation Setup
Generator and Retriever. We utilize two retrievers in our evaluations: GTE (Li et al., 2023) and BGE (Zhang et al., 2023). For generators, we evaluate four different models: Llama 3.1 Instruct-8B (Dubey et al., 2024), Command-R-7Bhttps://huggingface.co/CohereForAI/c4ai-command-r7b-12-2024, Microsoft Phi-4 (Abdin et al., 2024), and Gemma-2-2B (Team et al., 2024).
@@
Query Generation Setting. For IA, we employ few-shot prompting with GPT-4o to generate 30 queries based on the target document. We also use GPT-4o to generate a short description of the target document, summarizing its main idea and keywords. For details of different prompting strategies and the corresponding prompts for each stage, see Appendix B and Appendix F.
RAG Setting. As described in Section 2.1, we evaluate our attack in a more realistic setting compared to previous works, where the RAG system employs query-rewriting on the user’s query. We implement query-rewriting using a simple query-paraphrasing prompt via GPT-4o. We set $k=3$ for retrieval and investigate the impact of this hyperparameter across all attacks in Section 6.3.2. These retrieved documents are then provided as context to the generator via a system prompt. Details on both the query-paraphrasing and system prompts are presented in Appendix F. To demonstrate the impact of query-rewriting on inference, we also evaluate attacks in a vanilla RAG setup where query-rewriting is disabled (Appendix D).
Figure 4. ROC curves for Command-R (7B) as generator, GTE as retriever, across various datasets. Our attack (IA) achieves near-perfect inference across multiple datasets. ROC curves for other RAG configurations, can be found in Appendix H.
Baselines. We compare our attack with three prior black-box MIAs against RAG systems: RAG-MIA (Anderson et al., 2024), $S^{2}$ (Li et al., 2024c), and MBA (Liu et al., 2024b). RAG-MIA takes a simpler approach by directly probing the RAG system to ask if the target document appears in the context. $S^{2}$ uses the first half of the target document as a query and calculates the semantic similarity score (i.e., BLEU) between the RAG system’s responses and the original document as the membership score. They hypothesize that if a document is present in the database, the RAG system’s responses will exhibit high semantic similarity to the original content. MBA uses a proxy-LM to selectively mask words in the target document, followed by querying the RAG to predict those masked words. The number of successfully predicted masked words is used as the membership score. In our experiments, we use Qwen-2.5-1.5B (Yang et al., 2024) as the proxy LM.
@@ 6.2. Results
We posit that the superior performance of our attack stems from the utilization of multiple, diverse questions per document, in contrast to prior methods that typically rely on a single query. This multiplicity allows for a broader exploration of the document’s content, enhancing the likelihood of uncovering exploitable information (as detailed in Section 6.3.1). Similarly, the mask-based attack (MBA) benefits from analyzing multiple masked words within a single document, which might help explain its stronger performance relative to the other baselines.
@@ 6.3. Ablation Study
Here we evaluate the impact of varying several aspects of the RAG system and our attak.
@@ 6.3.2. Number of Retrieved Documents ($k$)
6.3.3. Distribution of Questions
Prior work shows that social and implicit biases can push LLMs toward generating affirmatively answered questions (Li et al., 2024a). To measure this tendency, we analyze the distribution of generated questions, ground-truth answers, and the outputs of the RAG system for these questions. We find a clear imbalance: for instance, with Gemma on TREC-COVID, questions answered with “yes” appear nearly 12 times more often than those answered with “no.” Accuracy is also skewed: 61% for ”yes” responses and 39% for “no”, suggesting a bias toward affirmative answers. To address this, we adapted the attack and scoring functions such that the model is given multiple answers to choose from, with only one being correct. Despite the format change, attack effectiveness remains comparable. For example, we sampled 100 members and 100 nonmembers from SCIDOCS and tested the attack against LLaMA 3.1 (without query rewriting). We found that attack performance remained almost unchanged: the AUC is 97.7% for yes/no questions and 97.8% for multiple-choice questions.
Figure 7. Attack performance (AUC) as a function of the UNK penalty $\lambda$. Performance increases steadily with higher $\lambda$ values before leveling off.
6.3.4. UNK Response Penalty ($\lambda$)
As described in (5), $\lambda$ is a hyperparameter that penalizes the RAG system when it cannot answer a question. We set $\lambda$ to a value greater than one (5) to reflect this intuition, but find that performance remains stable as long as it is reasonably large. For example, Gemma-2 on TREC-COVID shows an increase in attack AUC from 0.938 at $\lambda=0$ to 0.954 at $\lambda=7$, after which it stabilizes, as shown in Figure 7. This effect is further clarified by the proportion of UNK responses: 7.58% for members versus 54.05% for non-members when queried with our attack’s questions, highlighting the value of the penalty.
@@ 7.1. Assumptions on RAG Documents
Generic Documents. In addition to length, the targeted documents must contain information that the base LLM is unaware of or not very good at. The attack may perform poorly on highly generic documents, as they do not contain enough specific details to craft unique and distinguishable questions, and correspond to content that the base LLM may be familiar with. However, it is worth noting that in such cases, the utility of RAG might be limited, as generic documents provide less value for retrieval-based systems and the RAG owner might benefit in efficiency from discarding such documents from their datastore.
@@ 7.2. Analyzing Failure Cases
False Negatives. The fact that we observe high retrieval recall for our attack rules out the possibility of the target document being absent from the context provided to the RAG generator. The RAG’s inability to answer the question properly can thus have two potential reasons. On rare occasions, GPT-4o fails to paraphrase the user’s query accurately (see Appendix G for an example), which reflects a shortcoming in the RAG system—not being able to paraphrase a normal, benign query. For other cases, the RAG generator may struggle to answer the question even when the appropriate document is present in the provided context. Similarly, this can be attributed to the RAG’s generator lacking capabilities—especially given the fact that the question, by design, can be answered by GPT-4o-mini under the presence of the target document.
@@
False Positives. The RAG answering our queries correctly implies that the target document (corresponding to the query) is not required specifically as context to respond correctly. This can happen if similar documents with the necessary context are fetched by the retriever, or if the generator already possess sufficient knowledge to answer the question without relying on any context (see Appendix G for examples). To better understand this failure case, we compute the similarity between a non-member document $d$ and the document actually retrieved as context for a query corresponding to that non-member $d$, across multiple non-member documents and their corresponding queries generated for our attack. In Figure 9, we look at $n$-gram overlap and cosine similarity between retriever embeddings, and visualize them with respect to MIA scores for our attack. We observe that above some certain meaningful threshold ($0.2$ for 4-gram overlap, $0.9$ for embedding cosine similarity), there is a positive correlation between how ”similar” the non-member documents are to documents already present in the RAG, and the MIA Score (and by extension, questions answered correctly by the RAG). In summary, the failure cases are primarily due to limitations of the RAG system itself, such as occasional paraphrasing failures and the generator’s inability to answer questions effectively, rather than drawbacks of our attack.